OpenStudy (anonymous):

solve for x: log(5)(11-6x) = log(5)(1-x)

Nnesha (nnesha):

hint: \[\huge\rm log_b x = \log_b y\] \[\huge\rm\cancel { log_b} x = \cancel{\log_b} y\] x=y

OpenStudy (anonymous):

Ohhh... X = 2 right?

OpenStudy (anonymous):

Or is it no solution?

OpenStudy (pawanyadav):

Is 5 it's base

Nnesha (nnesha):

sorry i forgot :(

OpenStudy (anonymous):

Yes 5 is the base

OpenStudy (anonymous):

That's ok Nnesha. Thanks for trying to help :)

Nnesha (nnesha):

no i'm saying i forgot i ws helping u :P

OpenStudy (anonymous):

Oh hahaha

Nnesha (nnesha):

alright so what was your first step ?? :-)

OpenStudy (anonymous):

I would cross out the bases because they are the same right?

Nnesha (nnesha):

yep right and then simple algebra!!! :-)

OpenStudy (pawanyadav):

16-6x=1-x 15=5x x=?

Nnesha (nnesha):

and yes you got it right :-)

OpenStudy (mathmate):

Hint: dom log(x) = (0,inf)

OpenStudy (anonymous):

Oh this is a tricky no solution problem isn't it....

OpenStudy (mathmate):

... because?

OpenStudy (anonymous):

The base is 5 instead of 10?

OpenStudy (mathmate):

not exactly. Use the hint!

OpenStudy (anonymous):

Oh! I would have no 'y' value right?

Nnesha (nnesha):

nope when there is just log then u have to suppose 10 base

OpenStudy (mathmate):

actually, the hint applies to any base, so it was not indicated.

OpenStudy (anonymous):

Oh ok, so the answer would just be x = 2?

OpenStudy (mathmate):

want further hint? @lehmad

OpenStudy (anonymous):

yes please :p

Nnesha (nnesha):

i'm confused do you have to solve for x right ?

OpenStudy (mathmate):

Hint: Whenever you solve an equation involving log, you need to substitute back into the equation to reject all roots that make log negative (recall dom log(x)=(0,inf) ).

OpenStudy (mathmate):

@Nnesha yes, the solution x=2 is correct for the algebraic part, but not correct for the problem given!

OpenStudy (anonymous):

I have to solve for the problem given @Nnesha

OpenStudy (anonymous):

Oh ok, so when I plug x=2 into the final product, I get f(2) = (0,inf) which is not solvable so there is no solution

OpenStudy (mathmate):

\(log_5(11-6x)=log_5(-1)\) so x=2 is not an admissible solution because log(-1) is outside the domain of log.

OpenStudy (mathmate):

* -1 is outside the domain of log.

OpenStudy (mathmate):

sorry, gtg. be back later if you still have questions.

OpenStudy (anonymous):

I understand now! Thank you for bearing with me (haha) and helping me through all the steps!!!

OpenStudy (mathmate):

You're welcome! :)